### v2

# SBP-YOLO:A Lightweight Real-Time Model for Detecting Speed Bumps and Potholes

Abstract
Reliable and real-time detection of road speed bumps and potholes is crucial for anticipatory perception in advanced suspension systems, enabling timely and adaptive damping control. However, achieving high accuracy and efficiency on embedded platforms remains challenging due to limited computational resources and the small scale of distant targets. This paper presents SBP-YOLO, a lightweight and high-speed detection framework tailored for bump and pothole recognition. Based on YOLOv11n, the model integrates GhostConv and VoVGSCSPC modules into the backbone and neck to reduce computation while enhancing multi-scale semantic features. To improve small-object detection, a P2-level branch is introduced, with a Lightweight and efficient detection head (LEDH) mitigating the added computational overhead without compromising accuracy. A hybrid training strategy combining NWD loss, BCKD-based knowledge distillation, and Albumentations-driven augmentation further enhances localization precision and robustness. Experiments show that SBP-YOLO achieves 87.0% mAP, outperforming the YOLOv11n baseline by 5.8%. After TensorRT FP16 quantization, it runs at 139.5 FPS on Jetson AGX Xavier, delivering a 12.4% speedup over the P2-enhanced YOLOv11. These results validate the effectiveness of the proposed method for fast and low-latency road condition perception in embedded suspension control systems.
1 Introduction
Adjustable suspension systems, such as magnetorheological semi-active [1, 2] and fully active suspensions [3], offer real-time tunability of damping and stiffness to enhance ride comfort and handling stability [4]. However, their performance largely depends on timely and accurate perception of upcoming road conditions [5]. In the absence of reliable preview information, these systems are limited to reactive control, making it difficult to fully exploit their potential in suppressing body vibrations and improving vehicle dynamics. Among various road disturbances, speed bumps and potholes represent common yet critical anomalies that directly impact ride quality [6]. Early detection of such features enables predictive suspension control, allowing damping strategies to be adjusted in advance[7, 8, 9]. Existing methods for road anomaly detection rely on vehicle dynamics, vibration analysis, or multi-sensor fusion (e.g., accelerometers, gyroscopes, and GPS) [10, 11, 12]. While effective, these approaches are often task-specific, computationally intensive, or limited by high hardware costs, making them unsuitable for real-time embedded deployment in advanced suspension systems.
Camera-based visual perception offers a cost-effective and information-rich alternative to traditional sensor-based approaches for road condition monitoring. Deep learning has substantially advanced vision-based anomaly detection by enabling proactive identification of road defects such as speed bumps and potholes. Compared to acceleration signal analysis, vision-based methods provide earlier and more reliable detection, offering critical input for real-time suspension control. The YOLO (You Only Look Once) series has further improved the trade-off between detection accuracy and inference speed [13]. Bučko et al. [14] evaluated YOLOv3 for pothole detection under diverse lighting and weather conditions, and compared it with Sparse R-CNN. Their results showed that while Sparse R-CNN offered slightly higher robustness in extreme visual degradation, YOLOv3 remained competitive due to its lower hardware requirements and stable performance across most conditions. Subsequent versions introduced structural improvements: YOLOv4 adopted CSPDarknet and PANet for enhanced feature aggregation, while its lightweight variant YOLOv4-tiny was deployed by Park et al. [15] for pothole detection with superior real-time capability. ASAD et al. [16] implemented a lightweight YOLOv4-tiny model on a Raspberry Pi device for edge-based pothole detection, achieving a detection accuracy of 90% at 31.76 FPS. YOLOv5 [17] incorporated anchor-free heads and auto-learned anchor boxes, whereas YOLOv8 [18] leveraged decoupled heads and improved data augmentation techniques. Khan et al. [19] further optimized YOLOv8 with customized augmentations for autonomous driving scenarios. Building on these, Bhavana et al. [20] proposed POT-YOLO based on YOLOv8, achieving 97.6% accuracy in detecting complex pavement defects. More recently, YOLOv11 [21] introduced the C3k2 residual block and C2PSA attention mechanism to enhance multiscale feature representation. Its modular architecture spans from lightweight to large-scale variants, supporting flexible deployment across both embedded and high-performance platforms.
Despite progress in vision-based detection of potholes and speed bumps, real-world use under dynamic driving remains difficult. At long range, these anomalies appear small, making them hard to detect. Motion blur from fast driving degrades features, while shadows, glare, and dim lighting obscure road cues and reduce robustness. Limited computing power on embedded systems also limits input size and model depth, making real-time use harder. To address these challenges, this paper proposes a lightweight and efficient model tailored for speed bump and pothole detection. The main contributions are as follows:
  • $\bullet$
    A lightweight and efficiency detection head (LEDH) is introduced to reduce the computational overhead of the P2 layer, maintaining accuracy for small and distant targets. Combined with GhostConv modules and VoVGSCSPC blocks in the backbone and neck, the proposed architecture enhances feature extraction efficiency and robustness under diverse road and environmental conditions.
  • $\bullet$
    A comprehensive training strategy is adopted to improve detection performance in dynamic scenarios, incorporating NWD Loss [22] for precise localization, BCKD [23] for knowledge distillation, and Albumentations-based [24] data augmentation to simulate motion blur, lighting variations, and adverse weather conditions.
  • $\bullet$
    The model is quantized to FP16 and optimized using TensorRT, achieving an inference speed of 139.5 FPS on the NVIDIA Jetson AGX Xavier, thereby meeting real-time requirements for embedded road condition detection.
The remainder of this paper is organized as follows: Sect. 2 presents the proposed model architecture and the improved loss functions. Sect. 3 details the experimental setup and evaluation results. Sect. 4 discusses the model’s inference performance on the embedded platform. Finally, conclusions are drawn in Sect. 5.
2 Methods
2.1 SBP-YOLO Architecture
Figure 1: SBP-YOLO framework with Backbone, Neck, and Head. The Neck fuses multi-scale features (P2–P5) from the Backbone and passes them to the Head for classification and regression
This study focuses on the requirements of intelligent suspension systems, where accurate detection of road surface anomalies such as potholes and speed bumps is essential for predictive control. To address the constraints of limited computational resources and real-time inference, we adopt YOLOv11n as the baseline and introduce a series of structural improvements to form the proposed SBP YOLO architecture, as illustrated in Fig. 1.
The backbone network is enhanced by replacing conventional convolutional layers with GhostConv modules in deeper stages. This substitution significantly reduces computational cost while maintaining sufficient feature extraction capability. To improve the representation of features across different spatial resolutions, a new module named VoVGSCSPC is introduced. This design integrates densely connected patterns inspired by VoVNet with a GSConv-based bottleneck structure, and it is applied in both the backbone and the detection head to enhance semantic feature propagation.
To increase sensitivity to small and distant objects, an additional early-stage P2 detection layer is incorporated. This layer allows the model to capture fine spatial details by leveraging features from the shallow stages of the network. In contrast to the standard three-level detection structure (P3 to P5), SBP YOLO employs a four-level detection strategy that includes feature maps from P2 (160 $\times$ 160), P3 (80 $\times$ 80), P4 (40 $\times$ 40), and P5 (20 $\times$ 20). The detection head also integrates a lightweight and efficient design, referred to as LEDH, which further reduces computational overhead while preserving detection accuracy. All feature maps are annotated with their spatial dimensions, such as (320, 320, 16), where the three values represent the height, width, and number of channels, respectively. This notation provides a clear overview of the scale and structure of each layer.
Overall, SBP YOLO extends the baseline YOLOv11n by incorporating GhostConv modules, the VoVGSCSPC structure, a dedicated early detection branch, and the LEDH. These enhancements are designed to improve the model’s ability to capture fine spatial details, enhance feature representation across multiple scales, and reduce computational complexity. Such architectural improvements aim to enable accurate and efficient detection of road surface anomalies, supporting deployment in intelligent vehicle control and predictive suspension systems.
2.2 GhostConv Module
In practical road anomaly detection, identifying small and sparsely distributed targets—such as distant speed bumps and surface depressions—remains challenging due to their limited spatial coverage and low contrast with the surrounding environment. Effective detection of such anomalies requires a large receptive field while maintaining computational efficiency, which is particularly important for embedded platforms with constrained processing resources.
To address this issue, we incorporate the Ghost convolution (GhostConv) module from GhostNet [25] as a lightweight substitute for standard convolution (SC). GhostConv follows a two-stage strategy: an initial small-kernel convolution extracts primary features, which are then enriched by a depthwise convolution using a larger kernel (e.g., $5\times 5$) to generate supplementary "ghost" features. The depthwise convolution operates on half of the intermediate channels, allowing expansion of the receptive field with minimal computational overhead.
Integrating GhostConv into the backbone of SBP-YOLO enhances spatial feature representation while maintaining low complexity. This enables more accurate recognition of fine-grained road anomalies and supports efficient inference on low-power devices. As shown in Fig. 2, GhostConv significantly reduces the number of parameters and floating point operations (FLOPs) without sacrificing feature representation capability.
Let $c_{1}$ and $c_{2}$ represent the input and output channels, $k$ the kernel size, and $H\times W$ the spatial resolution. The parameter count and FLOPs for standard convolution are given by:
$P_{\text{SC}}=c_{1}c_{2}k^{2},\quad F_{\text{SC}}=2c_{1}c_{2}HWk^{2}.$ (1)
For GhostConv, the initial convolution uses kernel size $k_{m}$, and the depthwise convolution uses $k_{c}$. The corresponding parameter count and FLOPs are:
$P_{g}$ $=c_{1}\cdot\frac{c_{2}}{2}\cdot k_{m}^{2}+\frac{c_{2}}{2}\cdot k_{c}^{2},$ (2)
$F_{g}$ $=2c_{1}\cdot\frac{c_{2}}{2}\cdot HWk_{m}^{2}+2\cdot\frac{c_{2}}{2}\cdot HWk_{c}^{2}.$ (3)
The cost ratios relative to standard convolution are:
$\frac{P_{g}}{P_{\text{SC}}}=\frac{k_{m}^{2}+\frac{k_{c}^{2}}{c_{1}}}{2k^{2}},\quad\frac{F_{g}}{F_{\text{SC}}}=\frac{k_{m}^{2}+\frac{k_{c}^{2}}{c_{1}}}{k^{2}}.$ (4)
Under typical settings (e.g., $c_{1}\in[64,256]$, $k_{m}=3$, $k_{c}=5$), GhostConv reduces parameters and FLOPs to approximately 48%–55% of those of standard convolution. This demonstrates its effectiveness for efficient deployment in embedded detection frameworks. Moreover, Fig. 2(c) shows visualized feature maps from the 5th and 7th layers in the backbone. The GhostConv-enhanced model exhibits stronger responses to distant small-scale targets, highlighting its improved spatial awareness in complex road scenes.
Figure 2: Illustration of the Ghost module design and its impact on feature activation. (a) Ghost module; (b) GhostConv architecture; (c) Comparison of backbone activation heat maps at the 5th and 7th convolutional layers.
2.3 VoVGSCSPC Module
Figure 3: The flowchart of the VoVGSCSPC
Depthwise separable convolution (DSConv) reduces computational cost by decomposing standard convolution (SC) into a depthwise convolution (DWConv) for spatial filtering and a pointwise convolution (PWConv) for channel mixing. However, as spatial and channel operations are decoupled, DWConv suffers from limited inter-channel interaction, leading to feature isolation across groups. To address this, ShuffleNet [26] introduced channel shuffle to enable cross-channel information flow, while MobileNet [27] adopted extensive $1\times 1$ convolutions for dense channel fusion. Despite their effectiveness, these methods either suffer from reduced expressiveness or increased computational overhead, and often underperform SC in complex tasks.
GSConv [28] mitigates these drawbacks by combining the efficiency of GhostConv with a hardware-friendly channel shuffle mechanism. It enables effective intra- and inter-group feature communication, alleviating the isolation issue inherent in DWConv. By enhancing feature representation with minimal overhead, GSConv achieves accuracy comparable to or better than SC in lightweight models, making it well-suited for mobile and embedded visual applications.
As shown in Fig. 3, GSConv first applies a SC for downsampling, followed by a DWConv on the resulting features. The outputs are concatenated and shuffled to produce the final feature map:
$F_{\mathrm{GSC}}=\mathrm{Shuffle}\bigl{(}\mathrm{Cat}(\alpha(X_{C_{1}})_{C_{2}/2},\delta(\alpha(X_{C_{1}})_{C_{2}/2}))\bigr{)}_{C_{2}}$ (5)
where $X_{C_{1}}$ is the input with $C_{1}$ channels, $\alpha$ denotes SC, $\delta$ represents DWConv, and $F_{\mathrm{GSC}}$ indicate the output.
VoVGSCSPC [29] extends GSConv with a cross-stage architecture, splitting the input into two branches: one with a convolution followed by a GSBottleneck, and another with a DWConv residual branch. Their outputs are fused as
$GSB_{\mathrm{out}}=F_{\mathrm{GSC}}(F_{\mathrm{GSC}}(\alpha(X_{C_{1}})_{C_{1}/2}))+\alpha(X)_{C_{1}/2}$ (6)
$VoVGSCSP{C}_{\mathrm{out}}=\alpha\bigl{(}\mathrm{Concat}(GSB_{\mathrm{out}},\alpha(X_{C_{1}}))\bigr{)}$ (7)
where $GSB_{\mathrm{out}}$ and $VoVGSCSP{C}_{\mathrm{out}}$ denote intermediate and final outputs, respectively.
The VoVGSCSPC module, employed instead of the C3k2 modules in the backbone and neck, enhances multi-scale feature fusion and enriches semantic representation. This improvement enables the model to better capture both global and local context in images of potholes and speed bumps, thereby strengthening its perceptual and representational capabilities.
2.4 Lightweight and efficiency detection head (LEDH)
In suspension preview control, road anomalies such as speed bumps and potholes often appear as small, low-saliency objects at a distance, posing challenging small-object detection problems. Although YOLOv11 improves performance by introducing a shallow P2 detection layer, each detection head still employs two consecutive $3\times 3$ convolutions before the classification and regression branches, leading to considerable computational overhead and hindering deployment on embedded platforms.
To address this limitation, we propose the LEDH, a lightweight detection head that reduces redundancy while maintaining accuracy on small targets. As shown in Fig. 4, for each scale $i$, the input feature ${c}_{i}$ is first transformed by two stacked grouped convolutions (GroupConv) with group count $g=c_{i}/16$:
$F_{i}$ $=\mathrm{Conv}_{g}\big{(}\mathrm{Conv}_{g}({c}_{i})\big{)},$ (8)
$H_{i}$ $=\mathrm{Concat}\left[\mathrm{Conv}_{1\times 1}^{4r}(F_{i}),\ \mathrm{Conv}_{1\times 1}^{n_{c}}(F_{i})\right],$
where $F_{i}$ denotes the shared transformed feature map and $H_{i}$ the detection head output containing bounding box regression logits and class probabilities. Here, $n_{c}$ and $r$ are the number of classes and the Distribution Focal Loss (DFL) bin count, respectively.
Unlike YOLOv11, which applies separate convolutional blocks to the classification and regression branches, LEDH introduces a shared transformation stage with GroupConv, so that the enhanced feature representation is jointly utilized by both branches. This design eliminates redundant computation, significantly reducing parameters and FLOPs, while preserving spatial detail and semantic abstraction across scales. Subsequent lightweight $1\times 1$ convolutions produce the final predictions, yielding an efficient detection head well suited for real-time edge deployment.
Figure 4: Structure of the proposed LEDH. Multi-scale features (P2–P5) are transformed by two stacked $3\times 3$ GroupConv layers and then fed into parallel $1\times 1$ convolution branches for classification and regression.
2.5 Hybrid Loss Training Strategy
In road scene images, small structures such as speed bumps and potholes often appear with low pixel occupancy and blurred boundaries. This makes IoU-based losses highly sensitive to minor localization errors and insufficient for robust supervision. To mitigate this problem, we introduce the normalized Wasserstein distance (NWD) [22] as a distribution-aware metric, which models both predicted and ground truth boxes as 2D Gaussian distributions. NWD provides smoother gradients under positional shifts and better tolerance to scale variation. By combining NWD with conventional IoU loss, the model gains improved localization precision, especially for small objects. Given bounding boxes $A=(c{x_{a}},c{y_{a}},w_{a},h_{a})$ and $B=(c{x_{b}},c{y_{b}},w_{b},h_{b})$, the squared 2-Wasserstein distance between their corresponding Gaussian distributions $\mathcal{N}_{a}$ and $\mathcal{N}_{b}$ is computed. Since this distance does not directly serve as a similarity measure, it is normalized as:
$\mathrm{NWD}(\mathcal{N}_{a},\mathcal{N}_{b})=\exp\left(-\frac{\sqrt{W_{2}^{2}(\mathcal{N}_{a},\mathcal{N}_{b})}}{C}\right),$ (9)
where $C$ is a dataset-specific scaling constant (e.g., 0.5) determined via ablation studies. The final bounding box regression loss is defined as:
$\begin{cases}L_{\mathrm{NWD}}=1-\mathrm{NWD}(\mathcal{N}_{a},\mathcal{N}_{b})\\ L_{\mathrm{obj}}=(1-\alpha)\cdot L_{\mathrm{NWD}}+\alpha\cdot L_{\mathrm{iou}}\end{cases}$ (10)
where $L_{\mathrm{NWD}}$ represents the Wasserstein-based loss, $L_{\mathrm{iou}}$ is the standard IoU loss, and $\alpha=\mathrm{IoU}_{\mathrm{ratio}}\in[0,1]$ controls the loss weighting.
3 Experiments and Analysis
3.1 Experimental settings
The experiments were conducted on Ubuntu 24.04 with an Intel® Xeon E5-2680 CPU (2.4 GHz) and an NVIDIA Tesla V100 GPU (16 GB). The software environment comprised PyTorch 1.13.1 and CUDA 11.7. The model was trained using the Adam optimizer with an initial learning rate of 0.001, batch size of 16, and 300 epochs. Input images were resized to $640\times 640$ to enhance convergence and computational efficiency. Early stopping was employed, terminating training if validation performance did not improve for 30 consecutive epochs.
3.2 Dataset construction
Existing road surface datasets provide usable samples of potholes and speed bumps, but often lack comprehensive coverage of diverse driving conditions, such as varying lighting, occlusion, and road textures. To address this, we construct a targeted dataset by aggregating and refining relevant samples from multiple public sources, including benchmark datasets [30, 31, 14, 32] and community platforms such as Kaggle and Roboflow Universe. The resulting dataset captures representative road anomalies across complex real-world scenes, offering a more diverse and challenging benchmark for training and evaluating road perception models. Representative examples are shown in Fig. 5.
To further enhance generalization, data augmentation is applied to simulate challenging weather conditions such as rain and snow using efficient image transformation techniques [24]. Additional augmentations include random flipping and photometric adjustments (e.g., brightness, contrast, saturation), as illustrated in Fig. 6. The final dataset consists of 7,500 images, divided into 5,250 for training, 1,125 for validation, and 1,125 for testing.
Figure 5: Images of potholes and speed bumps in the dataset
Figure 6: Comparison of the effects after some image data augmentation
3.3 Evaluation indicators
To quantitatively assess detection performance, we adopt mean average precision ($\mathrm{mAP}$), including mAP at 50% IoU ($\mathrm{mAP}_{50}$) and the average over IoU thresholds from 50% to 95% ($\mathrm{mAP}_{50\text{-}95}$). Model efficiency is evaluated using the number of parameters (Params), GFLOPs, frames per second (FPS), and inference latency. Higher mAP reflects better detection accuracy, while higher FPS and lower latency indicate stronger real-time capability. GFLOPs and Params quantify computational and memory costs, which are critical for deployment on edge devices.
The evaluation metrics are defined as:
$P=\frac{TP}{TP+FP},\quad R=\frac{TP}{TP+FN}$ (11)
$\mathrm{AP}=\int_{0}^{1}P(R)\,dR,\quad\mathrm{mAP}=\frac{1}{M}\sum_{i=1}^{M}\mathrm{AP}_{i}$ (12)
where $TP$, FPFP, and $FN$ denote true positives, false positives, and false negatives, respectively. $P$ and $R$ represent precision and recall, and $M$ is the number of object categories.
3.4 Ablation experiment
Table 1: Ablation study by sequentially adding each module on top of the baseline. Each configuration includes all modules from the previous row plus the newly added component.
Model P (%) R (%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50\text{-}95}$ GFLOPs Params (M) FPSb=1
Baseline (YOLOv11n) 85.0 75.0 81.2 47.7 6.3 2.58 71.2
+ P2 Head 84.8 76.5 83.9 49.8 10.3 2.67 53.5
+ LEDH (replacing P2) 86.6 77.6 84.5 50.1 6.7 2.37 69.6
+ GhostConv 86.3 78.1 84.5 49.8 6.2 2.09 66.5
+ VoVGSCSPC 87.5 79.2 84.6 49.9 5.8 2.04 63.7
+ NWD Loss (Final) 89.0 79.4 86.6 51.1 5.8 2.04 63.7
Ablation studies in Table 1 evaluate the effectiveness of each proposed module by incrementally integrating them into the YOLOv11n baseline. Each configuration includes all modules from the previous row plus one newly added component.
The baseline YOLOv11n model demonstrates fast inference (71.2 FPS) but relatively limited detection accuracy ($\mathrm{mAP}_{50}=81.2\%$, $\mathrm{mAP}_{50\text{-}95}=47.7\%$), primarily due to the absence of low-level feature information critical for detecting small objects. Introducing a P2 detection head notably improves accuracy to $\mathrm{mAP}_{50}=83.9\%$, although this comes at the cost of increased computational complexity (FLOPs rise from 6.3G to 10.3G) and a drop in inference speed to 53.5 FPS.
To achieve a better accuracy-efficiency trade-off, we replace the standard P2 head with the proposed LEDH. This substitution yields further improvements in precision (+1.8%) and recall (+1.1%), while significantly reducing FLOPs by 35% and improving speed to 69.6 FPS.
Subsequently, the incorporation of GhostConv reduces the FLOPs by an additional 7.5% with minimal impact on detection accuracy. Building on this, the integration of the VoVGSCSPC backbone module not only reduces model complexity (to 5.8 GFLOPs and 2.04M Params), but also improves detection accuracy to $\mathrm{mAP}_{50}=84.6\%$.
Finally, the application of the NWD Loss further enhances the model’s ability to localize small and irregular objects, achieving the highest precision (89.0%) and overall detection performance ($\mathrm{mAP}_{50}=86.6\%$, $\mathrm{mAP}_{50\text{-}95}=51.1\%$) among all configurations. These results validate that each proposed component contributes to performance gains, and the final model presents a compelling balance of speed, accuracy, and lightweight design, making it well-suited for real-time road anomaly detection in resource-constrained environments.
3.5 Knowledge Distillation Experiment
Knowledge Distillation (KD) transfers knowledge from a larger, well-trained teacher model to a lightweight student model, aiming to enhance generalization with minimal computational overhead. The proposed KD framework adopts a hybrid strategy, integrating logit-level supervision using BCKD [23] and feature-level guidance via CWD [33]. Six corresponding intermediate layers are selected from both models to enable effective feature alignment. The detailed training hyperparameters are summarized in Table 2.
Table 2: Hyperparameter settings for the proposed Knowledge Distillation (KD) framework
Hyperparameter Value
Training Epochs 500
Batch Size 32
Optimizer SGD
Logits Distillation Loss BCKD
Feature Distillation Loss CWD
Distillation Loss Schedule Constant
Logits Loss Weight 0.5
Feature Loss Weight 0.5
Teacher Layers Used 12, 15, 18, 21, 24, 27
Student Layers Matched 12, 15, 18, 21, 24, 27
Table 3: Performance comparison of distilled model, teacher model, and non-distilled baseline on validation and test sets
Validation Set P (%) R (%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50\text{-}95}$
Baseline (Student) 89.0 79.4 86.6 51.1
+ Distillation 90.2 80.7 87.0 54.6
Teacher Model 91.0 81.4 87.4 53.8
Test Set P (%) R (%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50\text{-}95}$
Baseline (Student) 88.6 78.2 86.3 51.5
+ Distillation 89.9 81.8 87.7 54.9
Teacher Model 89.9 82.0 87.8 54.1
In our study, a teacher model with similar architecture and S-level scale (20.6 GFLOPs, 7.42M Params) was employed to guide the training of the student model. As shown in Table 3, KD leads to consistent improvements across all metrics. In particular, $\mathrm{mAP}_{50\text{-}95}$ increased by 3.5% on the validation set and 3.4% on the test set, indicating not only enhanced detection performance but also improved generalization. The comparable gains across both sets suggest that the student model benefits from the teacher’s guidance without exhibiting overfitting.
3.6 Model comparison experiment
To comprehensively evaluate detection quality and computational efficiency, we compare SBP-YOLO with representative YOLO variants (YOLOv5-YOLOv12) across standard performance and complexity metrics. As reported in Table 4, SBP-YOLO achieves an $\mathrm{mAP}_{50}$ of 86.6%, surpassing the YOLOv11n baseline by 5.4%, while reducing model parameters, FLOPs, and storage by 20.9%, 7.9%, and 17.9%, respectively. With knowledge distillation, SBP-YOLO (Distilled) further improves detection accuracy, reaching 87.0% $\mathrm{mAP}_{50}$ and 54.6% $\mathrm{mAP}_{50\text{-}95}$, without introducing additional computational overhead. These results highlight the effectiveness of our architecture and training strategy in achieving a strong trade-off between accuracy and efficiency.
Table 4: Comparison of SBP-YOLO with state-of-the-art YOLO models in terms of accuracy and efficiency
Models P (%) R (%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50\text{-}95}$ Params (M) GFLOPs Size (MB)
YOLOv5n [17] 83.4 74.5 80.6 46.1 2.18 5.8 4.43
YOLOv6n [34] 83.8 72.0 79.3 45.9 4.16 11.5 8.15
YOLOv8n [18] 88.2 73.5 82.4 48.3 2.68 6.8 5.36
YOLOv9t [35] 84.7 75.3 81.5 48.3 1.76 6.6 3.95
YOLOv10n [36] 84.8 72.5 82.0 48.3 2.27 6.5 5.48
YOLOv11n [21] 85.0 75.0 81.2 47.7 2.58 6.3 5.21
YOLOv12n [37] 84.1 74.2 80.9 47.5 2.56 6.3 5.25
SBP-YOLO (ours) 89.0 79.4 86.6 51.1 2.04 5.8 4.28
SBP-YOLO (distilled) 90.2 80.7 87.0 54.6 2.04 5.8 4.28
To further illustrate model performance during training, Fig. 7 presents the $\mathrm{mAP}_{50}$ and bounding box loss curves over 300 epochs. SBP-YOLO demonstrates consistently higher detection accuracy and lower training loss compared to other YOLO variants, indicating superior convergence speed and training stability. These trends further validate the effectiveness and robustness of SBP-YOLO for real-world deployment.
Figure 7: Training curves over 300 epochs for different YOLO models: (a) $\mathrm{mAP}_{50}$, (b) bounding box loss.
3.7 Performance Comparison
As shown in Fig. 8a, SBP-YOLO exhibits improved detection capability for small-scale road features at extended distances, such as far-field speed bumps and potholes. These targets often appear with reduced resolution and weak texture cues, posing challenges to existing detection models. Compared to baseline YOLO variants, SBP-YOLO achieves higher accuracy and greater consistency in identifying such long-range and low-visibility objects. Fig. 8b illustrates detection performance under varied illumination conditions, including sunset, nighttime, and intense sunlight. SBP-YOLO maintains stable accuracy across all scenarios. At night, it outperforms YOLOv5 by 9%, and during sunset—where reduced contrast can hinder detection—SBP-YOLO shows more robust results than other models. Under strong sunlight that induces glare, SBP-YOLO successfully identifies distant potholes, whereas models like YOLOv8 occasionally miss the target. To assess robustness under visually degraded environments, Fig. 8c reports results on augmented datasets simulating rain, snow, and motion blur. SBP-YOLO achieves 0.64 detection accuracy in simulated rain, outperforming YOLOv11n (0.47) and YOLOv5n (lowest). Similar improvements are observed under snow and motion blur conditions, demonstrating the model’s resilience across complex visual perturbations. These results demonstrate that SBP-YOLO provides a reliable and efficient solution for real-time detection of road surface features such as potholes and speed bumps. By identifying these geometric irregularities in advance, the model supports early adjustment of suspension parameters through feedforward control, helping to improve ride comfort and vehicle stability. This capability is particularly beneficial on complex or uneven roads, where timely perception is essential for safe and adaptive suspension behavior.
Figure 8: Qualitative comparison of SBP-YOLO and baseline models under challenging conditions: (a) detection of small distant road features; (b) robustness across varied lighting (sunset, night, strong sunlight); (c) performance under simulated rain, snow, and motion blur.
4 Experiment in embedded device
4.1 Experimental Environment
All embedded inference experiments were conducted on the NVIDIA Jetson AGX Xavier platform. As illustrated in Fig. 9, the experimental setup comprises the AGX Xavier module along with its carrier board. Inference was executed via the command line, enabling real-time monitoring of processing speed through the system terminal. The proposed SBP-YOLO model was converted and deployed using TensorRT, with the aid of an open-source optimization tool for YOLO models [38]. The software stack was based on NVIDIA JetPack 5.1.5.
Figure 9: Experimental hardware setup: (a) experimental platform; (b) display terminal for real-time inference results; (c) Jetson AGX Xavier-based hardware configuration .
4.2 Inference Experiments
To ensure a fair comparison, the baseline YOLOv11n model was augmented with a P2 detection head to align with the detection structure of SBP-YOLO. As a result, the performance evaluation focuses solely on the effectiveness of the proposed improvements in SBP-YOLO, rather than differences in model capacity or detection granularity. Table 5 and Fig. 10 report the inference performance across various precision modes, including INT8, FP16, and FP32. Under FP16, SBP-YOLO achieves a 15.39 FPS gain over the modified YOLOv11n; with FP32, the improvement is 13.52 FPS. Additionally, SBP-YOLO maintains a smaller model size across all configurations, reflecting its superior efficiency. When comparing different quantization modes under identical inference conditions, INT8 yields lower accuracy and frequent missed detections. FP16 maintains accuracy close to FP32, while significantly reducing latency. Notably, SBP-YOLO under FP16 consistently detects both long-range speed bumps and complex potholes, demonstrating a lower missed detection rate and higher reliability in real-time deployment scenarios.
Table 5: Inference time, frame rate, and model size comparison between SBP-YOLO and YOLOv11n (with P2 head) under different quantization settings.
Models Infer Time (ms) FPS Weights
YOLOv11n (INT8) 6.514 152.9 5.53 MB
YOLOv11n (FP16) 8.06 124.11 7.42 MB
YOLOv11n (FP32) 14.18 70.54 14.35 MB
SBP-YOLO (INT8) 6.42 155.78 5.24 MB
SBP-YOLO (FP16) 7.17 139.50 6.18 MB
SBP-YOLO (FP32) 11.9 84.06 9.63 MB
Figure 10: Comparison of recognition performance of different TensorRT quantization methods
5 Conclusion
This paper presented SBP-YOLO, a lightweight framework for real-time detection of speed bumps and potholes to support predictive suspension control. Building on YOLOv11, it incorporates a P2 detection head with LEDH employing shared GroupConv transformations to enhance small-object sensitivity while reducing computational overhead, alongside GhostConv and VoVGSCSPC modules for efficient multi-scale feature extraction. A hybrid training strategy combining NWD loss, backbone-level knowledge distillation, and advanced data augmentation improves robustness and generalization, achieving a 5.8% mAP gain over the YOLOv11n baseline. After TensorRT FP16 optimization, the model reaches 139.5 FPS on Jetson AGX Xavier, a 12.4% speedup over P2-enhanced YOLOv11n while maintaining superior accuracy. These results demonstrate the framework’s effectiveness for real-time embedded deployment, enabling anticipatory perception and adaptive damping control, with future work exploring infrared integration and multimodal detection to further enhance robustness under diverse conditions.